Rank each of the following value in order of smallest to largest: (7)

I. the upper quartile.

II. the median.

III. the 20th percentile.

IV. decile 6.

V. the 0.38 fractile.

VI. the 70th percentile.

VII. the lower quartile


1
Expert's answer
2021-05-07T14:24:38-0400

Let 'N' be the no. of bservation

then,


Upper Quartile, "Q_3=\dfrac{3N}{4}=0.75N"


Median, "Q_2=\dfrac{N}{2}=0.5N"


"20^{th}" ​Percentile,"P_{20}=\dfrac{20N}{100}=0.2N"


Decile 6, "D_6=\dfrac{6N}{10}=0.6N"


0.38 fractile, "F_{0.38}=0.38N"


"70^{th}" ​Percentile,"P_{70}=\dfrac{70N}{100}=0.70N"


Lower Quartile, "Q_1=\dfrac{N}{4}=0.25N"


From the above results, The ranks from smallest value to largest value are as follows-


(1) "20^{th}" Percentile

(2) Lower Quartile

(3) 0.38 fractile

(4) Median

(5) Decile 6

(6) "70^{th}" Percentile

(7) Upper Quartile